OpenStudy (anonymous):

1√8 how to rationalize the denominator

OpenStudy (dayakar):

is it\[\frac{ 1 }{ \sqrt{8} }\]

OpenStudy (anonymous):

Yes how do i rationalize the denominator

OpenStudy (dayakar):

\[\frac{ 1 }{ \sqrt{2*2*2} }\]

OpenStudy (dayakar):

\[=\frac{ 1 }{ 2\sqrt{2} }\]

OpenStudy (dayakar):

is it clear

OpenStudy (dayakar):

multiplying numerator and denominator with sqrt of 2

OpenStudy (dayakar):

\[=\frac{ 1 }{ 2sqrt{2} }*\frac{ \sqrt{2} }{ \sqrt{2}}\]

OpenStudy (dayakar):

\[=\frac{ \sqrt{2} }{ 2*2 }\]

OpenStudy (dayakar):

\[=\frac{ \sqrt{2} }{ 4 }\]

OpenStudy (dayakar):

rationalized the denominator